# Distributed Model Checking in Graphs Classes

### v1

@@ 5.1 The Case of Trees with Bounded Depth
For technical reasons that will appear clear later, let us introduce a weaker variant of the CONGEST model, called BROADCAST-CONGEST. This model is defined in the same way as the CONGEST model, except that, at each round, every node must communicate the same $\mathcal{O}(\log n)$-bit message to all its neighbors. That is, at every round, every node can solely broadcast a same $\mathcal{O}(\log n)$-bit message to its neighbors instead of potentially sending a specific $\mathcal{O}(\log n)$-bit message to each of its neighbors.
Lemma 15.
Let $\mathcal{A}$ be an algorithm such that, in every $n$-node network $G=(V,E)$, every node $u\in V$ is given an $\mathcal{O}(\log n)$-bit input $\textsf{In}(u)$, and must produce an $\mathcal{O}(\log n)$-bit output $\textsf{Out}(u)$. Let $R(n)$ be the round-complexity of $\mathcal{A}$ in the BROADCAST-CONGEST model. Then, there is a certification scheme using certificates on $\mathcal{O}(R(n)\log n)$ bits that certifies that a pair of functions $f:V\to\{0,1\}^{*}$ and $g:V\to\{0,1\}^{*}$ form an input-output pair for $\mathcal{A}$ (i.e., if, for every node $u\in V$, $\textsf{In}(u)=f(u)$, then, for every node $u\in V$, $\textsf{Out}(u)=g(u)$).
@@
Let us now prove a result analogous to Lemma 5, but for distributed certification.
Lemma 16.
Let $(G,L)$ be a connected $\Lambda$-labeled graph, and let $F\in\mathcal{F}_{d}$ be a subgraph of $G$. Let $\varphi(\overline{x})=\exists\overline{y}\;\zeta(\overline{x},\overline{y})$ be a formula in $\text{FO}[\Lambda]$, where $\zeta$ is an lca-reduced formula. Let us assume that every node $u$ of $G$ is provided with the following values:
@@
  • A set of labels $L(u)\subseteq\Lambda$, and a set of labels $\widehat{L}(u)\subseteq\hat{\Lambda}$;
@@
There is a certification scheme using $\mathcal{O}(\log n)$-bit certificates that certifies that $\textsf{true}(\varphi,F,L)=\textsf{true}(\widehat{\varphi},F,\tilde{L})$.
@@
Let $k=|\overline{x}|$. As in the proof of Lemma 4, let us assume w.l.o.g. that $k\geq 1$ (as otherwise we can create a dummy free variable $x$ for $\varphi$). Moreover, let us assume that $\widehat{\Lambda}$ and $\widehat{\varphi}$ are the set and the formula defined in the proof of Lemma 4. Since $\widehat{\Lambda}$ and $\widehat{\varphi}$ can be computed by every node without any communication, each node can reject in case they are not appropriately defined. Similarly, we can assume that $L$ and $\hat{L}$ are, respectively, a $\Lambda$-labeling and a $\hat{\Lambda}$-labeling. Thanks to Proposition 4, $\mathcal{O}(\log n)$-bits certificates are sufficient to certify that the encodings of $F$ and $T$ are correctly. The remaining of the proof consists in checking that $\widehat{L}$ is correctly set as specified the proof of Lemma 4.
Let us first suppose that $|\overline{y}|=1$. We shall deal with the case when $|y|>1$ later in the proof. We start by proving the result for the case where $\zeta$ is a lca-type $\textsf{type}_{\gamma,\delta}\in\textsf{Type}(k+1,d,\Lambda)$, where the variable $y$ is identified with variable $k+1$ in the definition of $\gamma$ and $\delta$. In a second stage, we’ll define $h$, $h_{x}$ and $h_{y}$ in the same way as we did in Lemma 4, and revisit the cases of Lemma 4.
In Case 1 and 2, observe that the algorithms provided in Lemma 5 are algorithms that fit with the broadcast constraint of the BCLIQUE model, running in $\mathcal{O}(d)$ rounds. On input $L$, these algorithms output $\widehat{L}$. By Lemma 15, we can certify whether $\widehat{L}$ is well defined in $d\cdot\mathcal{O}(\log n)=\mathcal{O}(\log n)$ rounds.
Now let us deal with Case 3. The algorithm $\mathcal{A}$ given in Lemma 5 for this case has three phases. The first two phases can be implemented by an algorithm $\mathcal{B}$ under the BCLIQUE model, running in $\mathcal{O}(d)$ rounds. The output of $\mathcal{B}$ consists in a marking of all active roots, as well as a marking of all good nodes. We can simulate these two first phases by giving a certificate to each node, interpreted as the markings given by $\mathcal{B}$, as well as a certificate of the execution of $\mathcal{B}$. According to Lemma 15, this can be implemented using certificates on $\mathcal{O}(\log n)$ bits. The last phase of Algorithm $\mathcal{A}$ aims at providing each node with the number of active roots. Each node $u$ receives a certificate consisting of:
@@
$\ell(\nu)=\begin{cases}\nu&\textsf{ if }\nu\leq k,\\ \infty&\textsf{ otherwise.}\end{cases}$
@@
    If $u\in F$ is marked as good by $\mathcal{B}$, then $\widehat{L}(u)=L(u)\cup\{\textsf{good},\ell(\rho(u))\}$;
@@
    If $u\in F$ is not marked as good by $\mathcal{B}$, then $\widehat{L}(u)=L(u)\cup\{\ell(\rho(u))\}$;
@@
    If $u\notin F$ then $\widehat{L}(u)=L(u)$.
We now check completeness and soundness of this certification scheme. Condition (1) is satisfied if and only if every node received the same value $\rho$ for the number of active roots. Conditions (2) and (3) are both satisfied if and only if $\textsf{val}(u)$ correspond to the number of active roots in the subgraph of $T$ rooted at $u$. In particular $\textsf{val}(r)$ is equal to the number of active roots. Then, conditions (1) to (4) are satisfied if and only if $\rho$ corresponds to the number of active roots in $F$. Finally, conditions (1) to (7) are satisfied if and only if $\widehat{L}$ is well defined according to the proof of Lemma 4.
Let us denote by $\mathcal{A}^{\textsf{type}}$ the verification algorithm, and by $\textsf{Cert}^{\zeta}$ the certificates described for each of the three cases when $\zeta$ is an lca-type formula. Let us now assume that $\zeta$ is an arbitrary lca-reduced formula. From Lemma 3, there exists a set $I\subseteq~{}\textsf{Type}(k+1,D,\Lambda)$ such that,
@@
For each $\psi\in I$, let us denote by $\widehat{\Lambda}^{\psi}$, $\widehat{\psi}$ and $\widehat{L}^{\psi}$ the reduction on $\mathcal{F}_{d}$ of $\Lambda$, $\psi$ and $L$, respectively. The prover provides each node $u$ of $G$ with the certificate $c^{\psi}(u)=\widehat{L}^{\psi}(u)$ for each $\psi$, in addition to the certificate $\textsf{Cert}^{\psi}$.
The verifier proceeds as follows. Each node $u$ checks that it accepts when running algorithm $\mathcal{A}^{\textsf{type}}$ with certificates $c^{\psi}(u)$, for every $\psi\in I$. Then, each node $u$ relabels the sets $\widehat{L}^{\psi}$ to make them disjoint, as we did in the proof of Lemma 4), and checks whether $\widehat{L}(u)=\bigcup_{\psi\in I}\widehat{L}^{\psi}(u)$. The soundness and completeness of the whole scheme follows from the soundness and completeness of $\mathcal{A}^{\textsf{type}}$.
It remains to deal with the case where $q=|\overline{y}|>1$. Let us denote by $\mathcal{A}$ the verification algorithm that we described for the case when $|\overline{y}|=1$. The prover provides the nodes with a set of triplets $(\zeta_{i},L_{i},\textsf{Cert}_{i})_{i\in[q]}$ such that, for each $i\in[q]$, $\text{Cert}_{i}$ is a certificate used to check whether
$\textsf{true}(\exists y_{i},\zeta_{i}(\overline{x},y_{1},\dots,y_{i}),G,L_{i})=\textsf{true}(\zeta_{i-1}(\overline{x},y_{1},\dots,y_{i-1}),G,L_{i-1}),$
Where $\zeta_{q}=\zeta$, $L_{q}=L$, $\zeta_{0}=\widehat{\varphi}$, and $L_{0}=\widehat{L}$. Every node $u$ accepts if $\mathcal{A}$ accept all certificates at $u$. From the soundness and completeness of $\mathcal{A}$, we obtain that every node accepts if and only if
$\textsf{true}(\varphi(\overline{x}),G,L)$ $=\textsf{true}(\exists(y_{1},\dots,y_{q})\,\zeta(\overline{x},(y_{1},\dots,y_{q})),G,L)$
$=\textsf{true}(\exists(y_{1},\dots,y_{q})\,\zeta_{q}(\overline{x},(y_{1},\dots,y_{q})),G,L_{q})$
$=\textsf{true}(\exists(y_{1},\dots,y_{q-1})\,\zeta_{q-1}(\overline{x},(y_{1},\dots,y_{q-1})),G,L_{q-1})$
@@
$=\textsf{true}(\exists y_{1}\,\zeta_{1}(\overline{x},y_{1}),G,L_{1})$
$=\textsf{true}(\zeta_{0}(\overline{x}),G,L_{0})$
$=\textsf{true}(\widehat{\varphi}(\overline{x}),G,\widehat{L}).$
@@
5.2 Certifying Treedepth Colorings and Skeletons
[22] have shown that, for every MSO property $\mathcal{P}$, there is a certification scheme using $\mathcal{O}(\log n)$-bit certificates for certifying $\mathcal{P}$ on graphs of bounded treedepth. Since the mere property “treedepth at most $k$” is expressible in MSO, we obtain the following.
Proposition 5 ([22]).
@@ 5.2 Certifying Treedepth Colorings and Skeletons
We show how to use this result for certifying a $(p,f(p))$-treedepth coloring. Moreover, we also extend this certification scheme to an encoding of a $p$-skeleton. Given a $(p,f(p))$-treedepth coloring $C_{p}$, and $p$-skeleton $\tau_{p}$, the encoding of $\tau_{p}$ is defined as the encoding of all decomposition forests $F^{U}$ of $G[U]$, for all $U\in\textsf{Col}(p)$. More precisely, in the encoding of $\tau_{p}$, for every $U\in\textsf{Col}(p)$, every node $u$ has a triple $\{(\textsf{root}^{U}(u),\textsf{parent}^{U}(u),\textsf{depth}^{U}(u))\}$ corresponding to the subtree $T_{u}^{U}$ containing $u$ of the decomposition forest $F^{U}$ of $G[U]$.
Lemma 17.
Let $f:\mathbb{N}\to\mathbb{N}$, and let $\mathcal{G}$ be a class of graphs of expansion $f$. Let $p>0$, and let us suppose that each node $u$ is given (1) a value $C_{p}(u)\in[f(p)]$, and (2) an encoding of $\tau_{p}$. There is a certification scheme using $\mathcal{O}(\log n)$-bit certificates for certifying that $C_{p}$ is a $(p,f(p))$-treedepth coloring of $G$, and that the encoding of $\tau_{p}$ is indeed defining a $p$-skeleton of $G$.
@@
Let us first certify that $C_{p}$ is a $(p,f(p))$-treedepth coloring of $G$. For every set $U\in\textsf{Col}(p)$, every node $u$ receives as a certificate: (1) The encoding of a rooted spanning tree $T$ of the component of $G[U]$ containing $u$ (where each component is identified with the node identifier of the root of $T$); (2) The certification of the property “treedepth at most $p$” for all the nodes in the component of $G[U]$ containing $u$. The soundness and completeness of (1) and (2) follow from Proposition 4 and Proposition 5.
Next, let us certify that $\tau_{p}$ is a $p$-skeleton of $G$. First, we use the verifier of Proposition 4 to check that the encoding of each tree $T_{u}^{U}$ is correct, for all $u\in V$ and all $U\in\textsf{Col}(p)$. If no nodes reject at that point, we can safely deduce that if $u$ and $v$ are in the same connected component of $G[U]$, then the encoding of $T_{u}^{U}$ and of $T_{v}^{U}$ must induce the same tree.
It remains to certify that each $T^{U}_{u}$ forms a decomposition tree of the component of $G[U]$ containing $u$. More precisely, we must certify that all neighbors of $u$ in $G[U]$ are either descendants or ancestors of $u$ in $T_{u}^{U}$. Let us fix an arbitrary set $U\in\textsf{Col}(p)$. Let $(\textsf{root}^{U}(u),\textsf{parent}^{U}(u),\textsf{depth}^{U}(u))$ be the encoding of $T_{u}^{U}$ that has $u$ in its input (given in the encoding of $\tau_{p}$). Then node $u$ receives the set $P^{U}_{u}=\{w^{u}_{1},\dots,w^{u}_{d}\}$ as certificate, corresponding to a sequence of at most $d\leq 2^{p}$ node identifiers representing the path from $\textsf{root}(u)$ to $u$. Then node $u$ accepts $P^{U}_{u}$ if the following holds:

### v2

@@ 6.1 Certifying Reduction of Trees with Bounded Depth
Lemma 6.2.
Let $\mathcal{A}$ be an algorithm such that, in every $n$-node network $G=(V,E)$, every node $u\in V$ is given an $\mathcal{O}(\log n)$-bit input $\textsf{In}(u)$, and must produce an $\mathcal{O}(\log n)$-bit output $\textsf{Out}(u)$. Let $R(n)$ be the round-complexity of $\mathcal{A}$ in the BROADCAST CONGEST model. Then, there is a certification scheme using certificates on $\mathcal{O}(R(n)\log n)$ bits that certifies that a pair of functions $f:V\to\{0,1\}^{*}$ and $g:V\to\{0,1\}^{*}$ form an input-output pair for $\mathcal{A}$ (i.e., if, for every node $u\in V$, $\textsf{In}(u)=f(u)$, then, for every node $u\in V$, $\textsf{Out}(u)=g(u)$).
@@
Let us now prove a result analogous to Lemma 4.1, but for distributed certification.
Lemma 6.3.
Let $(G,\ell)$ be a connected $\Lambda$-labeled graph, and let $F\in\mathcal{F}_{d}$ be a subgraph of $G$. Let $\varphi(\overline{x})=\exists\overline{y}\;\zeta(\overline{x},\overline{y})$ be a formula in $\mathrm{FO}[\Lambda]$, where $\zeta$ is an lca-reduced formula. Let us assume that every node $u$ of $G$ is provided with the following values:
@@
  • A set of labels $\ell(u)\subseteq\Lambda$, and a set of labels ^$\widehat{\ell}(u)\subseteq\hat{\Lambda}$;
@@
There is a certification scheme using $\mathcal{O}(\log n)$-bit certificates that certifies that $\textsf{true}(\varphi,F,\ell)=\textsf{true}(\widehat{\varphi},F,\tilde{\ell})$.
@@
Let $k=|\overline{x}|$. As in the proof of Lemma 3.7, let us assume w.l.o.g. that $k\geq 1$ (as otherwise we can create a dummy free variable $x$ for $\varphi$). Moreover, let us assume that $\widehat{\Lambda}$ and $\widehat{\varphi}$ are the set and the formula defined in the proof of Lemma 3.7. Since $\widehat{\Lambda}$ and $\widehat{\varphi}$ can be computed by every node without any communication, each node can reject in case they are not appropriately defined. Similarly, we can assume that $\ell$ and $\hat{\ell}$ are, respectively, a $\Lambda$-labeling and a $\hat{\Lambda}$-labeling. Thanks to Proposition 6.1, $\mathcal{O}(\log n)$-bits certificates are sufficient to certify that the encodings of $F$ and $T$ are correctly. The remaining of the proof consists in checking that $\widehat{\ell}$ is correctly set as specified the proof of Lemma 3.7.
Let us first suppose that $|\overline{y}|=1$. We shall deal with the case when $|y|>1$ later in the proof. We start by proving the result for the case where $\zeta$ is a lca-type $\textsf{type}_{\gamma,\delta}\in\textsf{Type}(k+1,d,\Lambda)$, where the variable $y$ is identified with variable $k+1$ in the definition of $\gamma$ and $\delta$. In a second stage, we’ll define $s$, $h$, $h_{s}$ and $h_{y}$ in the same way as we did in Lemma 3.7, and revisit the cases of Lemma 3.7. asdasdasd In Case 1 and 2, observe that the algorithms provided in Lemma 3.9 are algorithms that fit with the broadcast constraint of the BROADCAST CONGEST model, running in $\mathcal{O}(d)$ rounds. On input $\ell$, these algorithms output $\widehat{\ell}$. By Lemma 6.2, we can certify whether $\widehat{\ell}$ is well defined in $d\cdot\mathcal{O}(\log n)=\mathcal{O}(\log n)$ rounds.
Now let us deal with Case 3. The algorithm $\mathcal{A}$ given in Lemma 4.1 for this case has three phases. The first two phases can be implemented by an algorithm $\mathcal{B}$ under the BROADCAST CONGEST model, running in $\mathcal{O}(d)$ rounds. The output of $\mathcal{B}$ consists in a marking of all active roots, as well as a marking of all good nodes. We can simulate these two first phases by giving a certificate to each node, interpreted as the markings given by $\mathcal{B}$, as well as a certificate of the execution of $\mathcal{B}$. According to Lemma 6.2, this can be implemented using certificates on $\mathcal{O}(\log n)$ bits. The last phase of Algorithm $\mathcal{A}$ aims at providing each node with the number of active roots. Each node $u$ receives a certificate consisting of:
@@
$h(\nu)=\begin{cases}\nu&\textsf{ if }\nu\leq k,\\ \infty&\textsf{ otherwise.}\end{cases}$
@@
    If $u\in F$ is marked as good by $\mathcal{B}$, then ^$\widehat{\ell}(u)=\ell(u)\cup\{\textsf{good},h(\rho(u))\}$;
@@
    If $u\in F$ is not marked as good by $\mathcal{B}$, then ^$\widehat{\ell}(u)=\ell(u)\cup\{h(\rho(u))\}$;
@@
    If $u\notin F$ then ^$\widehat{\ell}(u)=\ell(u)$.
We now check completeness and soundness of this certification scheme. Condition (1) is satisfied if and only if every node received the same value $\rho$ for the number of active roots. Conditions (2) and (3) are both satisfied if and only if $\textsf{val}(u)$ correspond to the number of active roots in the subgraph of $T$ rooted at $u$. In particular $\textsf{val}(r)$ is equal to the number of active roots. Then, conditions (1) to (4) are satisfied if and only if $\rho$ corresponds to the number of active roots in $F$. Finally, conditions (1) to (7) are satisfied if and only if $\widehat{\ell}$ is well defined according to the proof of Lemma 3.7.
Let us denote by $\mathcal{A}^{\textsf{type}}$ the verification algorithm, and by $\textsf{Cert}^{\zeta}$ the certificates described for each of the three cases when $\zeta$ is an lca-type formula. Let us now assume that $\zeta$ is an arbitrary lca-reduced formula. From Lemma 3.4, there exists a set $I\subseteq~{}\textsf{Type}(k+1,D,\Lambda)$ such that,
@@
For each $\psi\in I$, let us denote by $\widehat{\Lambda}^{\psi}$, $\widehat{\psi}$ and ^$\widehat{\ell}^{\psi}$ the reduction on $\mathcal{F}_{d}$ of $\Lambda$, $\psi$ and $\ell$, respectively. The prover provides each node $u$ of $G$ with the certificate $c^{\psi}(u)=\widehat{\ell}^{\psi}(u)$ for each $\psi$, in addition to the certificate $\textsf{Cert}^{\psi}$.
The verifier proceeds as follows. Each node $u$ checks that it accepts when running algorithm $\mathcal{A}^{\textsf{type}}$ with certificates $c^{\psi}(u)$, for every $\psi\in I$. Then, each node $u$ relabels the sets ^$\widehat{\ell}^{\psi}$ to make them disjoint, as we did in the proof of Lemma 3.7), and checks whether ^$\widehat{\ell}(u)=\bigcup_{\psi\in I}\widehat{\ell}^{\psi}(u)$. The soundness and completeness of the whole scheme follows from the soundness and completeness of $\mathcal{A}^{\textsf{type}}$.
It remains to deal with the case where $q=|\overline{y}|>1$. Let us denote by $\mathcal{A}$ the verification algorithm that we described for the case when $|\overline{y}|=1$. The prover provides the nodes with a set of triplets $(\zeta_{i},\ell_{i},\textsf{Cert}_{i})_{i\in[q]}$ such that, for each $i\in[q]$, $\text{Cert}_{i}$ is a certificate used to check whether
$\textsf{true}(\exists y_{i},\zeta_{i}(\overline{x},y_{1},\dots,y_{i}),G,\ell_{i})=\textsf{true}(\zeta_{i-1}(\overline{x},y_{1},\dots,y_{i-1}),G,\ell_{i-1}),$
Where $\zeta_{q}=\zeta$, $\ell_{q}=\ell$, $\zeta_{0}=\widehat{\varphi}$, and $\ell_{0}=\widehat{\ell}$. Every node $u$ accepts if $\mathcal{A}$ accept all certificates at $u$. From the soundness and completeness of $\mathcal{A}$, we obtain that every node accepts if and only if
$\textsf{true}(\varphi(\overline{x}),G,\ell)$ $=\textsf{true}(\exists(y_{1},\dots,y_{q})\,\zeta(\overline{x},(y_{1},\dots,y_{q})),G,\ell)$
$=\textsf{true}(\exists(y_{1},\dots,y_{q})\,\zeta_{q}(\overline{x},(y_{1},\dots,y_{q})),G,\ell_{q})$
$=\textsf{true}(\exists(y_{1},\dots,y_{q-1})\,\zeta_{q-1}(\overline{x},(y_{1},\dots,y_{q-1})),G,\ell_{q-1})$
@@
$=\textsf{true}(\exists y_{1}\,\zeta_{1}(\overline{x},y_{1}),G,\ell_{1})$
$=\textsf{true}(\zeta_{0}(\overline{x}),G,\ell_{0})$
$=\textsf{true}(\widehat{\varphi}(\overline{x}),G,\widehat{\ell}).$
@@
6.2 Certifying Low-Treedepth Colorings and Skeletons
Bousquet et al. [22] have shown that, for every MSO property $\mathcal{P}$, there is a certification scheme using $\mathcal{O}(\log n)$-bit certificates for certifying $\mathcal{P}$ on graphs of bounded treedepth. Since the mere property “treedepth at most $k$” is expressible in MSO, we obtain the following.
Proposition 6.4 ([22]).
@@ 6.2 Certifying Low-Treedepth Colorings and Skeletons
We show how to use this result for certifying a $(p,f(p))$-treedepth coloring. Moreover, we also extend this certification scheme to an encoding of a $p$-skeleton. Given a $(p,f(p))$-treedepth coloring $C_{p}$, and $p$-skeleton $S$, the encoding of $S$ is defined as the encoding of all decomposition forests $F^{U}$ of $G[U]$, for all $U\in\textsf{Col}(p)$. More precisely, in the encoding of $S$, for every $U\in\textsf{Col}(p)$, every node $u$ has a triple $\{(\textsf{root}^{U}(u),\textsf{parent}^{U}(u),\textsf{depth}^{U}(u))\}$ corresponding to the subtree $T_{u}^{U}$ containing $u$ of the decomposition forest $F^{U}$ of $G[U]$.
Lemma 6.5.
Let $f:\mathbb{N}\to\mathbb{N}$, and let $\mathcal{G}$ be a class of graphs of expansion $f$. Let $p>0$, and let us suppose that each node $u$ is given (1) a value $C_{p}(u)\in[f(p)]$, and (2) an encoding of $S$. There is a certification scheme using $\mathcal{O}(\log n)$-bit certificates for certifying that $C_{p}$ is a $(p,f(p))$-treedepth coloring of $G$, and that the encoding of $S$ is indeed defining a $p$-skeleton of $G$.
@@
Let us first certify that $C_{p}$ is a $(p,f(p))$-treedepth coloring of $G$. For every set $U\in\textsf{Col}(p)$, every node $u$ receives as a certificate: (1) The encoding of a rooted spanning tree $T$ of the component of $G[U]$ containing $u$ (where each component is identified with the node identifier of the root of $T$); (2) The certification of the property “treedepth at most $p$” for all the nodes in the component of $G[U]$ containing $u$. The soundness and completeness of (1) and (2) follow from Proposition 6.1 and Proposition 6.4.
Next, let us certify that $S$ is a $p$-skeleton of $G$. First, we use the verifier of Proposition 6.1 to check that the encoding of each tree $T_{u}^{U}$ is correct, for all $u\in V$ and all $U\in\textsf{Col}(p)$. If no nodes reject at that point, we can safely deduce that if $u$ and $v$ are in the same connected component of $G[U]$, then the encoding of $T_{u}^{U}$ and of $T_{v}^{U}$ must induce the same tree.
It remains to certify that each $T^{U}_{u}$ forms a decomposition tree of the component of $G[U]$ containing $u$. More precisely, we must certify that all neighbors of $u$ in $G[U]$ are either descendants or ancestors of $u$ in $T_{u}^{U}$. Let us fix an arbitrary set $U\in\textsf{Col}(p)$. Let $(\textsf{root}^{U}(u),\textsf{parent}^{U}(u),\textsf{depth}^{U}(u))$ be the encoding of $T_{u}^{U}$ that has $u$ in its input (given in the encoding of $S$). Then node $u$ receives the set $P^{U}_{u}=\{w^{u}_{1},\dots,w^{u}_{d}\}$ as certificate, corresponding to a sequence of at most $d\leq 2^{p}$ node identifiers representing the path from $\textsf{root}(u)$ to $u$. Then node $u$ accepts $P^{U}_{u}$ if the following holds: